OpenStudy (anonymous):

A test consists of multiple choice questions worth 4 points and essay questions worth 8 points. The test has 22 questions that are worth a total of 100 points. Use a model to determine how many multiple choice questions are on the test. a. 14 b. 16 c. 19 d. 22

OpenStudy (anonymous):

x+y=22 4x+8y=100 now solve

OpenStudy (anonymous):

put x = number of essay questions and then solve \[8x+4(22-x)=100\]

OpenStudy (anonymous):

\[8x+88-4x=100\] \[4x+88=100\] \[4x=12\] \[x=3\]

OpenStudy (anonymous):

whats the answer ?

OpenStudy (anonymous):

3 essay, the rest multiple choice